17% of the households in a community donate to the local public television station. Among households that donate to the public television station, 30% donate to the local public radio station. Among households that do not donate to the public television station, only 5% donate to the public radio station. Answer the following for a randomly selected household in the community”

What is the probability a household donates to the public television station if the household donates to the public radio station?

Group of answer choices

0.449

0.051

0.567

0.7

0.551
Answer #1

Solution:

From the given information, we have summarized table as below:

Radio

Television

donate

Do not donate

Total

Donate

0.17*0.30 = 0.051

0.119

0.17

Do not Donate

0.83*0.05 = 0.0415

0.7885

0.83

total

0.0925

0.9075

1

We have to find the probability a household donates to the public television station if the household donates to the public radio station.

Required probability = 0.17*0.30 = 0.051

Answer: 0.051